Date Reported: 01/26/2015
Condition: Coffman will begin practice (weather permitting) on feb 2 on Fld #4. New Prospect is getting me their practice and game schedule ASAP. They will practice on Fld #3. Pitching rubbers need to be at 43' for both fields at this time.
Corrected? | Reported by: Pam Kemp
Date Corrected: | Time Corrected: | Corrected By:
Condition: Beginning Feb 2 Leoma Baseball team will begin practice. Bases need to be put down and pitching rubber. We are suppose to get a practice schedule from Leoma. I will fax when we receive it. (Beth spoke with Jerry 1-22-15)
Condition: Request from Jason Fox and LCHS soccer to paint lines mid Feb for soccer.
Date Reported: 01/26/2017
Condition: Swings and walking bridge(ask me) need links taken out to meet standards.
Corrected? YES | Reported by: Pam Kemp
Date Corrected: 1/27/2017 | Time Corrected: 11:00 am | Corrected By: Danny, Bobby, Chad
Action Taken: Raised swings to meet the standard
Date Reported: 01/27/2015
Condition: Roof is leaking on the police side of the Old Armory. Call John (629-1978)when you need to get in an look. Thanks
Corrected? | Reported by: John Sanders to Beth Campbell
Date Reported: 01/28/2015
Condition: Cold water supply line busted. Cut off water at road (meter on south side of impound lot). This feeds an old sink that is not used any more. Needs to be capped off. Can access room through door off gym floor. Call Beth for specifics. Thanks
Corrected? YES | Reported by: John Sanders to Beth Campbell
Date Corrected: 1/30/2015 | Time Corrected: 3:00 pm | Corrected By: Danny, Chad
Action Taken: Caped water line & replaced 5 windows
Date Reported: 01/30/2015
Condition: No hot water in office bldg.
Date Corrected: 1/30/2015 | Time Corrected: 1:00 p.m | Corrected By: Danny, Chad
Action Taken: Reset water heater.
Condition: Night lights on West end of Parks & Recreation bldg are not working
Date Corrected: 2/2/15 | Time Corrected: 1:00 p.m | Corrected By: Daren, Todd
Action Taken: Replaced bulb
Date Reported: 01/31/2017
Condition: Citizen called and reported Center light is not working
Date Corrected: 2/2/2017 | Time Corrected: 10:45 am | Corrected By: Jerry W Smith
Action Taken: No socket in middle light, there were four lights that work
Condition: 55 Gallon Drum full of used motor oil is in the shop. Need help with disposing of please.
Date Corrected: 2/2/2017 | Time Corrected: 9:00 a.m. | Corrected By: Jerry W Smith
Action Taken: Going to check with Farmers Supply